the harris and keller families decide to go to the movies together. the harris family lives 3 km west and 2 km north of the movie theater. the keller family lives 4 km east 3 km south of the movie theater. how far apart do the families live?
Answer by Jstrasner(112)

Hey,
So for this question, we need to use the Pythagorean theorem which is:
c^2 = a^2 + b^2 where c is the hypotenuse, or the longest side.
For this problem, since the houses are on complete opposite sides of the theater, then we need to use the Pythagorean theorem to find the distance the theater is from each house.
For the Harris house, a=2, b=3, so 2^2 + 3^2 = c^2 => 4+9 = 13 => 13=c^2
so the distance is the square root of 13 = 3.6
For the Keller house, a=3, b=4, so 3^2 + 4^2 = c^2 => 9+16=25 => 25=c^2
so the distance is the square root of 25 which is 5
Therefore:
The distance between the houses is 5 + 3.6 = 8.6
I hope this helps!